Let f(x) = 1/3x3 + x2 - 15x - 9. Use detailed sign tables in answering the following questions

(a) Find the intervals in which f is increasing or decreasing.

(b) Find the intervals in which the graph of y = f(x) is concave upwards or downwards.


Air is escaping from a spherical balloon at the rate of 2 cm3 per minute. How fast is the radius shrinking when the volume is 36 pi cm3.


Find the rate of change of the area A, of the circle with respect to its circumference C i.e dA/dC.

1
Expert's answer
2021-04-13T06:39:50-0400

a) Increase:


"(-\infty,-5)\ and\ (3,\infty)"

Decrease:


"(-5,3)"

b) Downwards:


"(-\infty,-1)"

Upwards:


"(-1,\infty)"